Temperature (X)

Ice Cream Sales (Y)

X2

Y2

XY

55

16

3025

256

880

60

22

3600

484

1320

65

25

4225

625

1625

72

30

5184

900

2160

75

37

5625

1369

2775

83

45

6889

2025

3735

85

59

7225

3481

5015

92

76

8464

5776

6992

95

89

9025

7921

8455

100

95

10000

9025

9500

Total

782

494

63262

31862

42457

There is a well-known correlation between temperature and ice cream sales. The hotter it gets, the more ice cream people eat! A local ice cream shop hires you to predict how busy they will be on a future date given the forecast. Using temperature (X) and number of ice cream cones sold during the lunch hour (Y), fill in the blank spaces in following chart to assess this temperature / ice cream cone sales relationship:

What is the correct regression intercept (a)? (round to three decimal places here)

            A). 782.000

            B). 0.708

            C). -2.983

            D). 553.656

            E). -92.432
Answer #1
X Y XY
total sum 782 494 42457 63262 31862

sample size ,n    =10   

here,   x̅   =Σx/n =78.2000,ȳ = Σy/n =49.4

  

SSxx = Σx² - (Σx)²/n =2109.600

SSxy=Σxy - (Σx*Σy)/n =3826.200

SSyy = Σy²-(Σy)²/n =7458.400

estimated slope , ß1 = SSxy/SSxx =3826.2/2109.6=1.8137

  

intercept,ß0 = y̅-ß1* x̄ = 49.4 - 1.8137*78.2 = -92.432

so, answer is option E)